OpenStudy (anonymous):

A 600kg truck travelling north at 5.0m/s collides with 4000kg truck moving west at 15m/s. If the two trucks remain locked together after impact, with what speed and in what direction do they move immediately after the collision?

OpenStudy (anonymous):

Hi there again Kristine... In order for you to develop this problem you need to devide it in two, X-axix and Y-axix. We know that when there is an "Inelastic" collision the formula is m1v1 + m2v2 = v(m1+m2) So: X-Axis: (0)(600) + (-15)(4000) = v(4000 + 600) note: The "minus" fifteen is the direction of the second object (west). V = -13.04 m/s >>> Again the "minus" is the direction (west) Y-Axis (5)(600) + (0)(4000) = v(600 + 4000) V= 0.652 m/s V in vector mode V = -13.04i + 0.652j(m/s) Let me know if this is helpful to u *-*

OpenStudy (anonymous):

thanks for your help :)